A house has incoming current I = 20 Amps. The standard 12 gauge copper wire in the house has a diameter of 0.0808 inches. What is the drift velocity of the incoming current? (Answer in m/s)

Vd =?
Answer #1

$Drift velocity $V_d=\frac{I}{neA}$ ; where I,n,e,A are current , electron density, electron charge, cross section area respectively.$ $ Given I=20A and $ e=1.6\times 10^{-19}C , n=8.4\times 10^{28}m^{-3} $ and $A=\pi \times 0.001026^{2} m^2. $ then $V_d=\frac{20}{8.4\times 10^{28}\times1.6\times 10^{-19}\times 3.308\times 10^{-6}}=4.5\times 10^{-4}m/s